# Kendall’s $\tau$ Coefficient for Logits Distillation

### v1

@@ Abstract
Knowledge distillation typically employs the Kullback-Leibler (KL) divergence to constrain the student model’s output to match the soft labels provided by the teacher model exactly. However, sometimes the optimization direction of the KL divergence loss is not always aligned with the task loss, where a smaller KL divergence could lead to erroneous predictions that diverge from the soft labels. This limitation often results in suboptimal optimization for the student. Moreover, even under temperature scaling, the KL divergence loss function tends to overly focus on the larger-valued channels in the logits, disregarding the rich inter-class information provided by the multitude of smaller-valued channels. This hard constraint proves too challenging for lightweight students, hindering further knowledge distillation. To address this issue, we propose a plug-and-play ranking loss based on Kendall’s $\tau$ coefficient, called Rank-Kendall Knowledge Distillation (RKKD). RKKD balances the attention to smaller-valued channels by constraining the order of channel values in student logits, providing more inter-class relational information. The rank constraint on the top-valued channels helps avoid suboptimal traps during optimization. We also discuss different differentiable forms of Kendall’s $\tau$ coefficient and demonstrate that the proposed ranking loss function shares a consistent optimization objective with the KL divergence. Extensive experiments on the CIFAR-100 and ImageNet datasets show that our RKKD can enhance the performance of various knowledge distillation baselines and offer broad improvements across multiple teacher-student architecture combinations.
@@ 1 Introduction
The recent advancements in deep neural networks (DNN) have significantly enhanced performance in the field of computer vision. This progress is driven by the pursuit of higher network capabilities, leading to an ongoing deepening of visual models and a continuous increase in the number of parameters. However, the heightened computational and storage costs associated with complex networks have limited their applicability. To address this, knowledge distillation (KD) has been proposed to obtain performant lightweight models. Typically, knowledge distillation involves a well-trained heavy teacher model and an untrained lightweight student model. The same data is fed to both the teacher and the student, with the teacher’s outputs serving as soft labels for training the student. By constraining the student to produce predictions that match the soft labels, the knowledge in the teacher model is transferred to the lightweight student model.
Most logit-based KD methods adhere to the paradigm introduced by Hinton [12], employing the Kullback-Leibler (KL) divergence [17] to align the logits output by the student and teacher models. The optimization goal of KL divergence is to achieve identical outputs between the student and teacher, thereby transferring as much knowledge as possible from the teacher to the student. However, given the large and complex nature of the teacher model, matching the output of the teacher without additional assistance is too challenging a task for the lightweight student. Recent studies [2; 27] have attempted to reduce the difficulty of the student’s learning task, such as reducing the gap between students and teachers by logits normalization, yet methods to fully leverage the information contained within logits remain underexplored.
Figure 1: Logits Value Distributions. (a) The original logits output by teacher and student. (b) The logits of teacher and student after a softmax with $T=4$.
@@
We notice that the optimization direction of KL divergence sometimes conflicts with the task loss optimization direction. As illustrated in Fig. 2, compared to Student 2, Student 1 exhibits a smaller KL divergence with the teacher; however, Student 2 achieves the correct classification result consistent with the teacher, while Student 1 does not. Therefore, a smaller KL divergence does not always lead to correct classification outcomes. On the other hand, the KL divergence places significant emphasis on the large channels, often at the expense of the smaller ones. In practice, however, the smaller-valued channels in logits constitute a substantial proportion of the teacher and student outputs on the training set (as shown in Fig. 1(a)). Although a temperature parameter can be added to the softmax equation to control the sharpness of the logits distribution, we still find this imbalance exists after softmax (as shown in Fig. 1(b)). We posit that these smaller-valued channels are equally rich in teacher knowledge, such as inter-class relationships, and that effectively utilizing this information is beneficial for the student’s learning.
In this paper, we introduce a plug-and-play Ranking Loss based on Kendall’s $\tau$ Coefficient [15] to assist the optimization process of KL divergence, called Rank-Kendall Knowledge Distillation (RKKD). This rank loss equitably focuses on the consistency of the magnitude order of values within logits, supplementing the student’s optimization process with knowledge from the smaller-valued channels and information about inter-class relationships in predictions. Moreover, the correct classification predicted by the teacher invariably possesses the top $1$st logit value, and the ranking loss constrains the correct category to the foremost rank, aiding the KL divergence in correcting the optimization direction to align with task loss. Specifically, we first explore three differentiable forms of Kendall’s $\tau$ coefficient and design the ranking loss function based on these. Subsequently, we demonstrate that the optimal optimization objective of the proposed ranking loss encompasses that of KL divergence, indicating that RKKD can assist the optimization process. By adding the ranking loss as an auxiliary loss to the existing knowledge distillation baseline, it can enhance performance for various teacher-student pairings in distillation tasks, suggesting that RKKD can serve as a plug-and-play method to improve the performance of knowledge distillation approaches. In conclusion, our contributions are as follows:
  • To mitigate the issues arising from the overly direct nature of KL divergence, which leads to the disregard of smaller-valued channels and suboptimal optimization directions, we introduce a plug-and-play ranking loss function to assist the optimization process in knowledge distillation (Rank-Kendall Knowledge Distillation, RKKD). Drawing on Kendall’s $\tau$ coefficient, we explore and discuss three different versions of the ranking loss function to improve the efficacy of knowledge distillation.
  • We establish that the proposed ranking loss function aligns with the optimal optimization objectives of KL divergence, and we substantiate this alignment through visualizations and empirical evidence, demonstrating the benefits of the ranking loss function for the optimization of distillation tasks.
@@
  • Extensive experiments were conducted on a variety of teacher-student architectures using the CIFAR-100 and ImageNet datasets. Our findings confirm the widespread effectiveness of the proposed ranking loss in distillation tasks, and its role as a plug-and-play auxiliary function provides substantial support for the training of distillation tasks.
@@ 2 Related Works
Knowledge Distillation. Knowledge distillation, initially proposed by Hinton [12], serves as a method for model compression and acceleration, aiming to transfer knowledge from a heavy teacher model to a lightweight student model. By feeding the same samples, the teacher can produce soft labels, and training the student with these soft labels allows the transfer of knowledge from the teacher model to the student. Knowledge distillation tasks can be mainly divided into two categories: feature-based distillation [3; 40; 36; 8; 21] and logit-based distillation [12; 20; 42; 14; 27; 5; 34]. Feature-based distillation additionally utilizes the model’s intermediate features, providing more information to the student and enabling the student to learn at the feature level from the teacher, often resulting in better performance. Considering safety and privacy, the intermediate outputs of the model are often not obtainable; hence, logit-based methods that only use the model outputs for distillation offer better versatility and robustness. Our proposed method can act as a plug-and-play module added to logit-based methods, offering higher flexibility and further enhancing the performance of logit-based methods.
@@
Recent knowledge distillation methods have found that overly strict constraints can sometimes hinder the student’s transfer of knowledge from the teacher’s soft labels. For instance, [2] discovered that differences in feature sizes in feature-based methods could limit the student’s learning; while [27] found that using the same temperature for both teacher and student in logit-based methods could affect further improvements in student performance. To reduce the learning difficulty for the student, [27; 2] provided methods to ease the logit matching difficulty between student and teacher, yet we find that methods providing additional guidance to the student have not been fully explored.
@@
Ranking Loss in Knowledge Distillation. For knowledge distillation, ranking loss is a relaxed constraint that can provide rich inter-class information. It was first applied in the distillation of recommendation systems [23; 29; 6; 22; 35], [18] explored the application of ranking loss in object detection tasks, and [7] discussed the role of ranking in the distillation of language model tasks. However, the exploration of ranking loss in logit-based image classification task distillation is not yet comprehensive. We find that the KL divergence used for distillation in classification tasks is an overly strict constraint that tends to overlook information from smaller-valued channels and may lead to suboptimal results. Our proposed method leverages ranking loss to balance the model’s attention to larger and smaller-valued channels, while also using inter-class relationships to help the model avoid suboptimal outcomes.
@@ 3 Preliminary
Most logit-based distillation methods adhere to the original KD proposed by Hinton [12], which transfers knowledge by matching the logit outputs of the student and teacher. This setting is more generalizable, allowing for distillation solely through outputs when the internal structures of the student and teacher are invisible. For a given dataset $\mathcal{D}$, assuming there are $C$ categories and $N$ samples, we possess a teacher model $f_{t}$ and a student model $f_{s}$. For a given sample $\mathcal{I}\in\mathcal{D}$, we can obtain the outputs of the teacher and student model, denoted as $z^{t}=f_{t}(\mathcal{I}),z^{s}=f_{s}(\mathcal{I})$ where $z^{t},z^{s}\in\mathbb{R}^{1\times C}$. Through a softmax function with temperature, the outputs are processed into the prediction vectors $q_{t},q_{s}\in\mathbb{R}^{1\times C}$ finally:
$q^{t}_{i}=\frac{exp(z^{t}_{i}/\mathcal{T})}{\sum_{j=1}^{C}exp(z^{t}_{j}/\mathcal{T})}$ (1)
$q^{t}_{i}=\frac{exp(z^{s}_{i}/\mathcal{T})}{\sum_{j=1}^{C}exp(z^{s}_{j}/\mathcal{T})}$ (2)
@@
$\mathcal{L}_{KL}(q^{t},q^{s})=\sum_{i=1}^{C}q_{i}^{t}\cdot log(\frac{q_{i}^{t}}{q_{i}^{s}})$ (3)
@@ 4 Ranking Loss Based on Kendall’s $\tau$ Coefficient
In this section, we introduce a ranking loss function designed to constrain the order of magnitude of channel values in the logits output by the student model. With the aid of the ranking loss, the knowledge distillation task can balance the overemphasis on larger logit values and the neglect of smaller ones as measured by the KL divergence. Additionally, the ranking loss imposes a constraint on the leading channel value, which helps to correct the optimization direction and avoid suboptimal solutions. In Section4.1, we will introduce the ranking loss function and employ Kendall’s $\tau$ coefficient to compute the ordinal consistency between the teacher and student logits. Section 4.2 will discuss the differentiable form of Kendall’s $\tau$ coefficient and, based on this, design three distinct forms of ranking loss functions. In Section 4.3, we will demonstrate that the optimization objectives of Rank Loss and KL divergence are aligned, and that Rank Loss can eliminate some of the suboptimal aspects of KL divergence, benefiting the distillation process for the student.
@@ 4.1 Differentiable Kendall’s $\tau$ Coefficient
In order to make the loss function pay more attention to the information provided by smaller channels as well as to help correct the suboptimal problem in Figure 1, we introduce the ranking loss based on Kendall’s $\tau$ coefficient. By pairing each of the $C$ channels, we can obtain $\frac{C(C-1)}{2}$ pairs, whose Kendall’s $\tau$ coefficient is expressed as follows:
@@
$\tau=\frac{\sum_{i}\sum_{j}sgn(z^{t}_{i}-z^{t}_{j})\cdot sgn(z^{s}_{i}-z^{s}_{j})}{\frac{1}{2}C(C-1)}$ (5)
where $sgn()$ represents sign function. While we have quantified the ordinal similarity between the teacher and student logits, the aforementioned formula is non-smooth. To utilize the ordinal similarity for gradient computation, we approximate the sign function with the $tanh$ function to convert it into a differentiable form:
$\tau_{d}$ $=\frac{\sum_{i}\sum_{j}tanh(z^{t}_{i}-z^{t}_{j})\cdot tanh(z^{s}_{i}-z^{s}_{j})}{\frac{1}{2}C(C-1)}$ (6)
$=\frac{2}{C(C-1)}\cdot\sum_{i}\sum_{j}\left(1-\frac{2}{1+e^{(z^{t}_{i}-z^{t}_{j})\cdot k}}\right)\cdot\left(1-\frac{2}{1+e^{(z^{s}_{i}-z^{s}_{j})\cdot k}}\right)$
where $k$ is a parameter that controls the steepness of the function; a larger $k$ causes the $tanh$ function to more closely approximate the sign function. By negating the similarity measure in Eq. 6, it can serve as a loss function to enforce the consistency of the logits order between the teacher and student:
$\mathcal{L}_{RK}=-\tau_{d}=-\frac{2}{C(C-1)}\cdot\sum_{i}\sum_{j}\left(1-\frac{2}{1+e^{(z^{t}_{i}-z^{t}_{j})\cdot k}}\right)\cdot\left(1-\frac{2}{1+e^{(z^{s}_{i}-z^{s}_{j})\cdot k}}\right)$ (7)
@@
4.2 Different Forms of Ranking Loss
In the initial application of the ranking loss[43], it is necessary to compute the gradients for two input vectors separately. However, In the scenario of distillation, the soft labels from the teacher model do not require gradients. Therefore, we propose three variants of the diff-Kendall ranking loss suitable for distillation scenarios, aimed at further exploring the role of ranking loss in distillation. From Eq.5, we can derive an equivalent form:
$\tau=\frac{\sum_{i}\sum_{j}sgn((z^{t}_{i}-z^{t}_{j})\cdot(z^{s}_{i}-z^{s}_{j}))}{\frac{1}{2}C(C-1)}$ (9)
For Eq.5 and Eq.9, we can transform them into a differential form of ranking loss, and gradients are not computed for the output part corresponding to the teacher, which is:
$L_{\tau}^{form1}=\frac{\sum_{i}\sum_{j}tanh(z^{t}_{i}-z^{t}_{j})_{detach}\cdot tanh(z^{s}_{i}-z^{s}_{j})}{\frac{1}{2}C(C-1)}$ (10)
$L_{\tau}^{form2}=\frac{\sum_{i}\sum_{j}tanh[(z^{t}_{i}-z^{t}_{j})_{detach}\cdot(z^{s}_{i}-z^{s}_{j})]}{\frac{1}{2}C(C-1)}$ (11)
Further, since the gradient of the teacher’s output is not required, we can directly use the sign function instead of the tanh function to obtain the ranking loss:
$L_{\tau}^{form3}=\frac{\sum_{i}\sum_{j}sgn(z^{t}_{i}-z^{t}_{j})_{detach}\cdot tanh(z^{s}_{i}-z^{s}_{j})}{\frac{1}{2}C(C-1)}$ (12)
4.3 How RKKD Benefit Optimizing Logits Distillation

### v2

# Enhancing Logits Distillation with Plug&Play Kendall’s $\tau$ Ranking Loss

@@ Abstract
Knowledge distillation typically minimizes the Kullback–Leibler (KL) divergence between teacher and student logits. However, optimizing the KL divergence can be challenging for the student and often leads to sub-optimal solutions. We further show that gradients induced by KL divergence scale with the magnitude of the teacher logits, thereby diminishing updates on low-probability channels. This imbalance weakens the transfer of inter-class information and in turn limits the performance improvements achievable by the student. To mitigate this issue, we propose a plug-and-play auxiliary ranking loss based on Kendall’s $\tau$ coefficient that can be seamlessly integrated into any logit-based distillation framework. It supplies inter-class relational information while rebalancing gradients toward low-probability channels. We demonstrate that the proposed ranking loss is largely invariant to channel scaling and optimizes an objective aligned with that of KL divergence, making it a natural complement rather than a replacement. Extensive experiments on CIFAR-100, ImageNet, and COCO datasets, as well as various CNN and ViT teacher-student architecture combinations, demonstrate that our plug-and-play ranking loss consistently boosts the performance of multiple distillation baselines. Code is available at https://github.com/OvernighTea/RankingLoss-KD
Machine Learning, ICML
@@ 1 Introduction
The recent advancements in deep neural networks (DNN) have significantly enhanced performance in the field of computer vision. However, the heightened computational and storage costs associated with complex networks limite their applicability. To address this, knowledge distillation (KD) has been proposed to obtain performant lightweight models. Typically, knowledge distillation involves a well-trained heavy teacher model and an untrained lightweight student model. The same data is fed to both the teacher and the student, with the teacher’s outputs serving as soft labels for training the student. By constraining the student to produce predictions that match the soft labels, the knowledge in the teacher model is transferred to the lightweight student model.
@@
Most logit-based KD methods adhere to the paradigm introduced by Hinton (Hinton et al., 2015), employing the Kullback-Leibler (KL) divergence (Kullback & Leibler, 1951) to align the logits output by the student and teacher models:
$\mathcal{L}_{KL}(q^{t}\,\|\,q^{s})=\sum_{i=1}^{C}q_{i}^{t}\cdot\mathrm{log}(\frac{q_{i}^{t}}{q_{i}^{s}}).$ (1)
Where $q_{t},q_{s}\in\mathbb{R}^{1\times C}$ is the prediction vectors of teacher and student. The optimization goal of KL divergence is to achieve identical outputs between the student and teacher, thereby transferring as much knowledge as possible from the teacher to the student. However, the optimization process of KL divergence is not easy, as it is prone to suboptimal points, which can hinder further improvement in student performance. As illustrated in Fig. 1, compared to Student 2, Student 1 exhibits a smaller KL divergence with the teacher; however, Student 2 achieves the correct classification result consistent with the teacher, while Student 1 does not. This indicates that the optimization direction of KL divergence sometimes diverges from the task objective, leading students to suboptimal points.
Intuitively, and as a preliminary observation, Eq. 1 shows that in the KL divergence each channel is weighted by the predicted probability of the teacher. Channels with small $q_{i}^{t}$ therefore contribute marginally to the loss, causing the student to tend to largely ignore them (the formal derivation is provided in Appendix A.5). In practice, most channels in a single prediction have lower probabilities, as illustrated in Fig. 2. Neglecting these channels hampers knowledge transfer for two reasons:
R1: Gradient attenuation. Because low-probability channels carry tiny weights, the corresponding gradients are strongly damped, which makes alignment with the teacher network difficult..
R2: Impaired capture of inter-class information. Low-probability channels retain rich cross-category correlation information in the teacher logits; overlooking them deprives the student of these semantic correlations. For example, in Fig. 1, Student 1 may not learn the distinction between the cat and fox classes.
Figure 2: Most Channels in the Distribution are Low-Probability Channels.
Considering these challenges, we aim to find an auxiliary method that mitigates the issues caused by channel scale differences, learns inter-class relationship information, and avoids the suboptimal points that KL divergence might encounter while maintaining the optimization objective of KL divergence. To address this problem, we propose a ranking-based plug-and-play auxiliary loss. The benefits of imposing ranking constraints are as follows:
B1: Gradients from the ranking term depend only weakly on the absolute channel scale.
B2: Ordered channel scores make cross-category relationships explicit.
B3: Fixing relative channel order helps the model escape suboptimal solutions.
Therefore, we propose to constrain the channel ranking similarity between student and teacher. We construct a plug-and-play ranking loss function based on Kendall’s $\tau$ Coefficient. This ranking loss can supplement the attention to low-probability channels in the logits and provide inter-class relationship information. We demonstrate that the gradients provided by KL divergence are related to channel scale, whereas the proposed ranking loss is not. We also show that the optimization objective of the proposed ranking loss is consistent with that of KL divergence, and visualize the loss landscape to illustrate that the ranking loss helps avoid suboptimal points in the early stage and does not alter the optimization objective in the end. Extensive experiments on CNN and ViT show that the proposed plug-and-play ranking loss can enhance the performance of various logit-based methods. In conclusion, our contributions are as follows:
  • We introduce a plug-and-play ranking loss for assisting knowledge distillation tasks, addressing the issues of KL divergence’s neglect of low-probability channels and its tendency to fall into suboptimal points, while also learning inter-class relationship information to further enhance student performance.
  • We demonstrate that the gradients provided by KL divergence exhibit sensitivity to channel scales, while the proposed ranking loss remains more robust to such variations. We also prove and visualize that the optimization objective of the proposed ranking loss is consistent with that of KL divergence and helps avoid the suboptimal points of KL divergence.
@@
  • Extensive experiments are conducted on a variety of CNN and ViT teacher-student architectures using the CIFAR-100, ImageNet, and MS-COCO datasets. Our findings confirm the widespread effectiveness of the proposed ranking loss in various distillation tasks, and its role as a plug-and-play auxiliary function provides substantial support for the training of distillation tasks.
@@ 2 Related Works
Knowledge Distillation. Knowledge distillation, initially proposed by Hinton (Hinton et al., 2015), serves as a method for model compression and acceleration, aiming to transfer knowledge from a heavy teacher model to a lightweight student model. By feeding the same samples, the teacher can produce soft labels, and training the student with these soft labels allows the transfer of knowledge from the teacher model to the student. Knowledge distillation tasks can be mainly divided into two categories: feature-based distillation (Chen et al., 2022; Zhang & Ma, 2021; Yang et al., 2022c; Guo et al., 2023; Park et al., 2019; Yang et al., 2022b) and logit-based distillation (Hinton et al., 2015; Li et al., 2023; Zhao et al., 2022; Jin et al., 2023; Sun et al., 2024; Chi et al., 2023; Wen et al., 2021; Yang et al., 2023b). Feature-based distillation additionally utilizes the model’s intermediate features, providing more information to the student and enabling the student to learn at the feature level from the teacher, often resulting in better performance. Considering safety and privacy, the intermediate outputs of the model are often not obtainable; hence, logit-based methods that only use the model outputs for distillation offer better versatility and robustness. Our proposed method can act as a plug-and-play module added to logit-based methods, offering higher flexibility and further enhancing the performance of logit-based methods.
@@
Recent knowledge distillation methods have found that overly strict constraints can sometimes hinder the student’s transfer of knowledge from the teacher’s soft labels. For instance, (Cao et al., 2022) discovered that differences in feature sizes in feature-based methods could limit the student’s learning; while (Sun et al., 2024) found that using the same temperature for both teacher and student in logit-based methods could affect further improvements in student performance. To reduce the learning difficulty for the student, (Sun et al., 2024; Cao et al., 2022) provided methods to ease the logit matching difficulty between student and teacher, yet we find that methods providing additional guidance to the student have not been fully explored.
@@
Ranking Loss in Knowledge Distillation. For knowledge distillation, ranking loss is a relaxed constraint that can provide rich inter-class information. It was first applied in the distillation of recommendation systems (Reddi et al., 2021; Tang & Wang, 2018; Choi et al., 2021; Qin et al., 2023; Yang et al., 2022a), (Li et al., 2022) explored the application of ranking loss in object detection tasks, and (Gao et al., 2020) discussed the role of ranking in the distillation of language model tasks. However, the exploration of ranking loss in logit-based image classification task distillation is not yet comprehensive. We find that the KL divergence used for distillation in classification tasks tends to overlook information from low-probability channels and may lead to suboptimal results. Our proposed method leverages ranking loss to balance the model’s attention between high-probability and low-probability channels, while also using inter-class relationships to help the model avoid suboptimal outcomes.
@@ 3 Preliminary
Most logit-based distillation methods adhere to the original KD proposed by Hinton (Hinton et al., 2015), which transfers knowledge by matching the logit outputs of the student and teacher. This setting is more generalizable, allowing for distillation solely through outputs when the internal structures of the student and teacher are invisible. For a given dataset $\mathcal{D}$, assuming there are $C$ categories and $N$ samples, we possess a teacher model $f_{t}$ and a student model $f_{s}$. For a given sample $\mathcal{I}\in\mathcal{D}$, we can obtain the outputs of the teacher and student model, denoted as $z^{t}=f_{t}(\mathcal{I}),z^{s}=f_{s}(\mathcal{I})$ where $z^{t},z^{s}\in\mathbb{R}^{1\times C}$.
Through a softmax function with temperature, the outputs are processed into the prediction vectors $q_{t},q_{s}\in\mathbb{R}^{1\times C}$ finally:
$q^{t}_{i}=\frac{\mathrm{exp}(z^{t}_{i}/\mathcal{T})}{\sum_{j=1}^{C}\mathrm{exp}(z^{t}_{j}/\mathcal{T})},\quad q^{t}_{i}=\frac{\mathrm{exp}(z^{s}_{i}/\mathcal{T})}{\sum_{j=1}^{C}\mathrm{exp}(z^{s}_{j}/\mathcal{T})}$ (2)
@@
$\mathcal{L}_{KL}(q^{t}\,\|\,q^{s})=\sum_{i=1}^{C}q_{i}^{t}\cdot\mathrm{log}(\frac{q_{i}^{t}}{q_{i}^{s}})$ (3)
@@ 4 Ranking Loss Based on Kendall’s $\tau$ Coefficient
In this section, we introduce the plug-and-play ranking loss function designed to constrain the ranking of channels in the logits output by the student model. With the aid of the ranking loss, the knowledge distillation task can balance the overemphasis on larger logit values and the neglect of smaller ones as measured by the KL divergence. Additionally, the ranking loss imposes a constraint on the leading channel value, which helps to correct the optimization direction and avoid suboptimal solutions. In Sec. 4.1, we will introduce the ranking loss function and employ Kendall’s $\tau$ coefficient to compute the ordinal consistency between the teacher and student logits. In Sec. 4.2, we discuss how ranking loss benefits optimizing logits distillation with KL divergence. Furthermore, we discuss the differentiable form of Kendall’s $\tau$ coefficient and, based on this, design three distinct forms of ranking loss functions in Appendix A.3.
@@ 4.1 Differentiable Kendall’s $\tau$ Coefficient
In order to make the loss function pay more attention to the information provided by low-probability channels as well as to help correct the suboptimal problem in Fig. 1, we introduce the ranking loss based on Kendall’s $\tau$ coefficient. By pairing each of the $C$ channels, we can obtain $\frac{C(C-1)}{2}$ pairs, whose Kendall’s $\tau$ coefficient is expressed as follows:
@@
$\tau=\frac{\sum_{i}\sum_{j<i}\mathrm{sgn}(z^{t}_{i}-z^{t}_{j})\cdot\mathrm{sgn}(z^{s}_{i}-z^{s}_{j})}{\frac{1}{2}C(C-1)}$ (5)
where $\mathrm{sgn}()$ represents sign function. While we have quantified the ordinal similarity between the teacher and student logits, the aforementioned formula is non-smooth. To utilize the ordinal similarity for gradient computation, we approximate the sign function with the $\tanh$ function to convert it into a differentiable form:
$\tau_{d}$ $=\frac{\sum_{i}\sum_{j<i}\tanh(k\cdot(z^{t}_{i}-z^{t}_{j}))\cdot\tanh(k\cdot(z^{s}_{i}-z^{s}_{j}))}{\frac{1}{2}C(C-1)}$ (6)
$=\frac{2}{C(C-1)}\cdot\sum_{i}\sum_{j<i}\left(1-\frac{2}{1+e^{2\cdot(z^{t}_{i}-z^{t}_{j})\cdot k}}\right)\cdot$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\left(1-\frac{2}{1+e^{2\cdot(z^{s}_{i}-z^{s}_{j})\cdot k}}\right)$
where $k$ is a parameter that controls the steepness of the function; a larger $k$ causes the $\tanh$ function to more closely approximate the sign function. By negating the similarity measure in Eq. 6, it can serve as a loss function (bounded within $[-1,1]$) to enforce the consistency of the logits order between the teacher and student:
$\mathcal{L}_{RK}=-\tau_{d}$ (7)
@@
4.2 How Ranking Loss Benefits Optimizing Logits Distillation with KL Divergence
4.2.1 From the Perspective of Gradient: Ranking Loss Cares about Low-Probability Channels
In this section, we analyze why ranking loss cares about low-probability channels. The gradient of KL divergence and ranking loss is written as follows. The specific calculation process of the gradient can be found in the appendix A.5:
$\frac{\partial\mathcal{L}_{\text{KD}}}{\partial z^{s}_{i}}=-T\left(q^{t}_{i}-q^{s}_{i}\right).$ (9)
$\frac{\partial\mathcal{L}_{RK}}{\partial z^{s}_{i}}=-\frac{k}{C(C-1)}\sum_{j\neq i}\left[1-\tanh^{2}\left(k(z^{s}_{i}-z^{s}_{j})\right)\right]\cdot$ (10)
$\tanh\left(k(z^{t}_{i}-z^{t}_{j})\right)$
Figure 3: The Gradient of Ranking Loss and KL Divergence for Different Values of Logits.
The gradient scale of the KL divergence is influenced by the scales of the teacher and student models. Consequently, for outputs from low-probability channels, when the scales of the teacher and student are similar, the KL divergence’s gradient for these components becomes negligible, leading to the neglect of information from low-probability channels. Although temperature scaling is typically employed to address this issue, it uniformly amplifies the gradients of all logits, thereby continuing to overlook the outputs of low-probability channels. In contrast, the gradients produced by ranking loss will be less affected by the scale of the teacher’s logits. The gradient of a logit channel primarily depends on the difference between its rank and the target rank, effectively harnessing the knowledge from low-probability channels. Therefore, In the early stage of training, ranking loss helps the model quickly learn the overall ranking of the teacher model, which helps the model converge to a better initial solution faster.
Table 1: CIFAR-100 Heterogeneous Architecture Results. The Top-1 Accuracy (%) is reported as the evaluation metric. The teacher and student have heterogeneous architectures. We incorporate the proposed ranking loss into the existing logit-based methods, with the performance gains indicated in parentheses. The best and second best results are emphasized in bold and underlined.
KD Teacher ResNet32$\times$4 ResNet32$\times$4 ResNet50 ResNet32$\times$4 WRN-40-2
79.42 79.42 79.34 79.42 75.61
Student WRN-16-2 WRN-40-2 MN-V2 SHN-V1 SHN-V1
73.26 75.61 64.60 70.50 70.50
FitNet (Adriana et al., 2015) 74.70 77.69 63.16 73.59 73.73
CRD (Tian et al., 2019) 75.65 78.15 69.11 75.11 76.05
ReviewKD (Chen et al., 2021) 76.11 78.96 69.89 77.45 77.14
DIST (Huang et al., 2022) 75.58 78.02 68.66 76.34 76.00
LSKD (Sun et al., 2024) 77.53 79.66 71.19 76.48 76.93
KD (Hinton et al., 2015) 74.9 77.7 67.35 74.07 74.83
KD+Ours 75.18(+0.28) 78.50(+0.80) 70.45(+3.10) 75.98(+1.91) 76.13(+1.30)
CTKD (Li et al., 2023) 74.57 77.66 68.67 74.48 75.61
CTKD+Ours 75.71(+1.14) 78.61(+0.95) 70.18(+1.51) 76.67(+2.19) 76.80(+1.19)
DKD (Zhao et al., 2022) 75.7 78.46 70.35 76.35 76.33
DKD+Ours 75.99(+0.29) 78.75(+0.29) 70.90(+0.55) 77.36(+1.01) 76.43(+0.10)
MLKD (Jin et al., 2023) 76.52 79.26 71.04 77.18 77.44
MLKD+Ours 76.83(+0.31) 79.86(+0.60) 71.66(+0.62) 77.63(+0.45) 77.87(+0.43)
In our study, we conduct a visualization of the gradients associated with both the ranking loss and the KL divergence loss in Fig .3. The logits of the student model are randomly generated but maintain the same scale as those of the teacher model across each channel. Our findings indicate that the ranking loss consistently provides gradients of similar scale across varying logit values, whereas the gradients from the KL loss are heavily dependent on the specific logit values. This suggests that the ranking loss offers a more uniform attention distribution across different logits compared to the KL loss. Notably, the ranking loss assigns low-probability gradients to logits with larger values, as they are the classification targets and reach the correct rank. Consequently, these logits receive less attention from the ranking loss.
4.2.2 From the Perspective of Optimal Solution Domain: Ranking Loss won’t Interfere with KL Divergence at the End